# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Results
Fig 7: Variability of estimated model parameters. A, B Distributions of estimated field and coupling parameters under the active and passive conditions for mouse $574078$. C: Scatter plots of coupling strength of reciprocal pairs under the active (red) and passive (blue) conditions for mouse $574078$. The coupling asymmetries were $0.147$ (active) and $0.105$ (passive). The asymmetry was assessed by the average absolute difference of the reciprocal couplings $\langle|\bar{\theta}_{ij}-\bar{\theta}_{ij}|\rangle_{ij}$, where $\bar{\theta}_{ij}$ indicates time-average of $\theta_{ij,t}$ and $\langle\cdot\rangle_{ij}$ refers to the average over the combinations of $i,j$. D-F Group-level (all mice) comparisons for the original dataset: field variance (D), coupling variance (E), and coupling asymmetry (F). G-I: Plots analogous to D-F for shuffle-subtracted parameter variances and coupling asymmetry. Each subplot of D-I contains the p-values of Wilcoxon signed-rank tests for the active vs. passive conditions.
Estimating time-dependent parameters of the kinetic Ising model for mice V1 data: population results
@@
We assessed key features in estimated models for all mice by comparing them with trial-shuffled data. Figure 7A-C illustrates these features using an exemplary mouse ($5571078$). Figure 7A, B shows distributions of time-averaged fields $\theta_{i,t}$ and couplings $\theta_{ij,t}$ under the active and passive conditions, while Figure 7C shows a scatter plot of time-averaged reciprocal couplings $\theta_{ij,t}$ vs $\theta_{ji,t}$ to evaluate coupling asymmetry. In the active condition, the medians of field parameters decreased, reflecting reduced firing rates, while the medians of couplings remained near zero in both conditions. Field and coupling parameter variances increased, and coupling asymmetry strengthened in the active condition. These trends were consistent across all mice (Fig. 7D-F). These features are key aspects of neural dynamics that are closely related to entropy flow. However, we also note that they are not entirely independent variables of each other, nor solely responsible for determining entropy flow.
@@
While increased parameter variabilities and coupling asymmetry were observed under the active condition, they may be influenced by the lower neuronal activity. To examine this, we compared results with trial-shuffled data across all mice. Figures 7G-I show field and coupling variances in both conditions, adjusted by subtracting shuffled data values for each mouse. Notably, observed values in both active and passive conditions were significantly higher than shuffled data ($p=2.91\times 10^{-11}$ (active), $p=1.103\times 10^{-7}$ (passive) for fields, $p=4.676\times 10^{-8}$ (active), $p=1.455\times 10^{-11}$ (passive) for couplings, Wilcoxon signed-rank test). Note that the observed significant heterogeneity in the field parameters is likely associated with the coupling heterogeneity. These results confirm that the variability observed in active or passive conditions are not explained by noise couplings. The coupling asymmetry was higher than shuffled results only for the active condition ($p=1.185\times 10^{-5}$(active) and $p=0.1287$ (passive) for asymmetry).
@@
Comparisons of these significant changes of the parameter variability (i.e., shuffled results subtracted) between the active and passive conditions showed significantly greater values in the active condition ($p=8.273\times 10^{-4}$ for fields, $p=6.421\times 10^{-4}$ for couplings, Wilcoxon signed-rank test, Fig.7G,H), indicating greater variabilities in both field and coupling parameters during active behavior. A similar analysis of the mean couplings across mice revealed slightly but significantly larger values under the active condition (Fig. S5 Fig). In contrast, coupling asymmetry showed no significant difference ($p=0.1287$, Wilcoxon signed-rank test, Fig.7I). The lack of change in asymmetry in the effective couplings validates the use of the proposed mean-field method for comparing the coupling effect, which primarily arises from variability changes rather than asymmetry. These findings validate enhanced parameter variability in the sparse neuronal activity during active engagement.
Entropy flow dynamics
@@
Fig 8: A Time courses of dissipative entropy for the original (solid) and shuffled data (dash-dot) under the active and passive conditions. The plot also shows time courses of the averaged population spike rates (dashed). The same color is used for the lines under the same condition. B Total entropy flows summed across all time bins for each mouse under the active and passive conditions. The same mouse is connected by a line. C Shuffled-subtracted total entropy flow (original - shuffle), plotted per mouse under the active and passive conditions.
@@
Using the estimated parameters of the state-space kinetic Ising model, we computed entropy flow dynamics. Figure 8A shows the time-varying entropy flow of a representative mouse ($574078$) under the active and passive conditions (red and blue solid lines, respectively). In both cases, transient increases in entropy flow coincided with declines in the mean population spike rate (dashed lines). Similar patterns appeared across all mice analyzed (S6 Fig). These increases align with the second law, indicating that greater entropy dissipation is required when the system is transitioning to a lower entropy state, characterized by reduced firing rates.
@@
The entropy flow time courses for this mouse showed no clear differences between the active and passive conditions. To assess population-level effects, we analyzed all 37 mice and computed total entropy flow across time bins for each condition (Fig. 8B). The comparison revealed significantly lower total entropy flow in the active condition ($p=0.01159$, Wilcoxon signed-rank test). Note that neurons exhibited reduced firing rates (Fig. S3 Fig) and increased parameter variability (Fig. 7D,E) during the active condition.
@@
To isolate the effect of couplings, we compared the observed total entropy flows with shuffled data results (Fig.8C). The estimated entropy flow for shuffled data includes the impact of firing rate dynamics and estimation error on couplings from other neurons; therefore, subtracting shuffling results from observed entropy flow isolates contributions of couplings among different neurons beyond the sampling fluctuation. Positive values of the shuffle-subtracted total entropy flow in both conditions indicate that the couplings caused a significant entropy flow increase ($p=1.455\times 10^{-11}$ for active, $p=1.455\times 10^{-11}$ for passive, Wilcoxon signed-rank test). These shuffle-subtracted entropy flows behave in agreement with the theoretical prediction by the Sherrington-Kirkpatrick model [37]. In the active condition, the increased coupling variability (and asymmetry) from the shuffle-subtracted values were positively correlated with the shuffle-subtracted entropy flows, while the increased field heterogeneity was negatively correlated (S7 Fig.A-C). These effects disappeared in the passive condition, possibly due to small changes in the variabilities and asymmetry introduced by shuffling (S7 Fig.D-F).
@@
We analyzed the differences in coupling-related entropy flows between the active and passive conditions for all mice (Fig.8C). The result shows no significant difference between the two conditions ($p=0.1448$, Wilcoxon signed-rank test). However, these coupling-related entropy flows having the same magnitude emerged under distinct neural activity states: sparser, lower activity with increased variability in field and coupling parameters in the active condition; less sparse, higher activity with reduced variability in the passive condition. Thus, coupled with the previous results, this result indicates that the greater coupling variability in the active condition led to increased total entropy flow, making it comparable to the passive condition despite significantly sparser firing rate distributions. Consistent with this view, we note that a recent study by Aguilera et al. using this dataset complementarily reported that a lower bound on entropy production, derived under steady-state assumptions using a variational framework, was higher in the active condition when normalized per spike [13].
Model-based perturbation analysis
@@
Fig 9: Model-based perturbation analysis A Entropy flow $\sigma_{t}$ of mouse 574078 in the active condition, computed after rescaling the fitted parameters as $\boldsymbol{\theta}\to\beta\boldsymbol{\theta}$. The dashed line indicates $\beta=1$. B Forward entropy flow, $\sigma_{t}^{\mathrm{forward}}$. C Reverse entropy flow, $\sigma_{t}^{\mathrm{backward}}$. D Shuffle-subtracted entropy flow in the active condition, isolating the contribution of interactions beyond firing rate dynamics. E Shuffle-subtracted entropy flow in the passive condition, showing interaction-driven peaks at $\beta<1$. F Comparison of shuffle-subtracted entropy flow between active and passive conditions across all mice, showing significantly higher values in the active condition (Wilcoxon test, $p=1.455\times 10^{-10}$). Shuffle-subtracted entropy flow is obtained over a low–gain range $\beta\in[0.2,1.0]$ and across all bins. Lines connect active (left) to passive (right) for each mouse.
@@
To further elucidate the difference in the estimated entropy flow in active and passive conditions, we performed a model-based perturbation analysis by rescaling the fitted model parameters as $\boldsymbol{\theta}\to\beta\boldsymbol{\theta}$ and computing the resulting entropy flow to assess its sensitivity to parameter perturbation. An example result from mouse $57407$ (Fig.9A) shows that the entropy flows during stimulus presentation and waiting (gray image) periods exhibited distinct behaviors in response to the scaling. The transient increases in entropy flow caused by firing rate reduction at stimulus onset and offset persisted as the scaling parameter $\beta$ increased. In contrast, we observed the entropy flow peaked at $\beta<1$ during the waiting period.
@@
Both forward and reverse conditional entropies ($\sigma_{t}^{\rm forward}$ and $\sigma_{t}^{\rm backward}$ in Eq. 12) decreased with increasing $\beta$ during the waiting period (Fig.9B,C), indicating that both processes became more deterministic. This trend suggests that, as $\beta$ increases, the system transitions from a disordered phase toward a ferromagnetic phase, rather than into a quasi-chaotic regime, as expected at a phase transition[43]. Thus, these results indicate that the subsampled neural population during this period does not operate in a quasi-deterministic chaotic regime, but rather in a subcritical regime.
@@
By subtracting the entropy flow estimated from trial-shuffled data, which preserves only firing rate dynamics, we confirmed that the two bands of increased entropy flow during stimulus presentation are attributable to firing rate changes, whereas the increase at $\beta<1$ arises from interactions, since the former disappeared but the latter persisted after shuffle subtraction (Fig.9D). Notably, the interaction-driven entropy flow during the waiting period peaked at $\beta<1$.
More interestingly, the interaction-driven entropy flow revealed by parameter scaling differed significantly between the active and passive conditions (Fig.9D,E), a result that was confirmed across all mice (Fig.9F). Notably, the previous analysis at $\beta=1$ showed no difference in shuffle-subtracted (i.e., interaction-driven) entropy flow between the two conditions (Fig.8C). Thus, the model-based perturbation analysis uncovered differences in entropy flow between active and passive states that were not apparent at $\beta=1$.
Relation between entropy flow and behavioral performance
@@
Fig 10: Correlations between neural dynamics and behavioral performance. A A histogram of firing-rate differences (active–passive) across neurons for mouse 574978 with an image im036_r. A skewness of this distribution was used as a sparsification index. B The skewness vs behavioral performance as measured by d-prime for 37 mice exposed to three images. C Mean-field rate as a function of time-average mean $g_{i}$ and variability $\Delta_{i}$ of their inputs for a mouse 574978. The color map shows theoretical mean-field rate under the steady-state (Eq. 19). D Mean-field entropy flow of individual neurons as a function of time-average mean and variability of their inputs. The color map shows theoretical entropy flow under the steady-state (Eq. 20). E Entropy flow normalized by activity rate vs behavioral performance. F Entropy flow normalized by activity rate obtained from trial-shuffled data vs behavioral performance. In E and F, there is one outlier mouse below -20, which was included in the statistical analysis.
@@
Finally, we investigated the relationship between neural dynamics and behavioral performance across individual mice. In the following analyses, we extended the analysis to include two additional images (im012_r and im115_r).
@@
First, we examined how sparseness, assessed from individual neurons’ activity rates, relates to behavioral outcomes. As shown in S3 Fig, neuronal activities were significantly reduced under active conditions, accompanied by increased sparsity of firing-rate distributions. To further characterize this effect, we examined whether the reduction was uniform across neurons or driven by a subset of neurons by computing the skewness of the firing-rate difference between active and passive conditions (Fig. 10A). A uniform reduction results in a skewness of zero, whereas negative skewness indicates that only some neurons decreased their activity, reflecting the sparsification. We found that this sparsification index was significantly correlated with behavioral performance, indicating that task engagement is reflected in changes in sparsity quantified at the level of individual neurons’ activity rates (Fig. 10B).
@@
Having established the link between activity‑rate sparsity and behavior, we next turned to entropy flow to ask whether it provides additional explanatory power beyond rate changes alone. The variability of effective couplings were significantly higher during the active condition. To gain insight into the contributions of couplings to entropy flow, we computed the activity rate and mean‑field entropy flow of individual neurons as a function of the mean and variability of their inputs. In the low‑input and stationary regime, entropy flow increased with both higher mean input and greater variability. We observed that neurons receiving high mean input tended to have less variable inputs, whereas neurons with low mean input exhibited larger variability (Fig. 10C,D). These results suggest that total entropy flow is shaped not only by high‑input (typically high‑firing) neurons but also by low‑input neurons with high variability.
@@
These patterns imply two separable sources of entropy flow: (i) mean‑input–driven contributions that track high firing, and (ii) variability‑driven contributions that can be substantial even at low firing. To more focus on the latter, we considered entropy flow per activity rate. This normalization reduces the direct dependence on mean rate and makes variability‑driven effects, particularly those arising in low‑rate neurons, observable on equal footing with high‑rate effects. The shift in mean entropy flow per activity rate of individual neurons (active - shuffle) was significantly correlated with behavioral performance (Fig. 10E). Moreover, this correlation was weaker and non‑significant for trial‑shuffled data, indicating contributions from highly variable couplings during active conditions (Fig. 10F). This finding suggests that the thermodynamic cost per spiking activity is related to mice performance, with couplings contributing in addition to activity‑rate sparsity.
@@
As an alternative explanation, behavioral performance could be related to entropy‑flow changes concentrated in high‑firing neurons. We therefore tested whether neurons with higher spike rates tended to increase entropy flow during active engagement in mice with higher task performance (Supporting Information 5 text). While this tendency correlated significantly with behavioral performance for one image, it was not significant for the other two images. We therefore infer that high‑firing–based changes alone cannot consistently account for performance differences. Instead, the more robust association with entropy flow per activity rate supports a complementary role of variability‑driven, coupling‑mediated contributions – including those from low‑rate neurons – in explaining behavioral performance.

### v3

@@ Results
Figure 7: Variability of estimated model parameters. A, B Distributions of the time-averaged field values $\bar{\theta}_{i}$ (A) and the time-averaged couplings $\bar{\theta}_{ij}$ (B) for mouse $574078$ under the active and passive conditions. The shaded violin plots depict kernel-density estimates of the empirical distributions; gray dots are the underlying observations (one dot per neuron in A, one per coupling in B). Short horizontal caps at the top and bottom indicate the sample maximum and minimum, respectively. Horizontal red bars mark the mean, while horizontal green bars mark the median. C: Scatter plots of coupling strength of reciprocal pairs under the active (red) and passive (blue) conditions for mouse $574078$. The coupling asymmetries were $0.147$ (active) and $0.105$ (passive). The asymmetry was assessed by the average absolute difference of the reciprocal couplings $\langle|\bar{\theta}_{ij}-\bar{\theta}_{ji}|\rangle_{ij}$, where $\bar{\theta}_{ij}$ indicates the time-average of $\theta_{ij,t}$ and $\langle\cdot\rangle_{ij}$ refers to the average over the combinations of $i,j$. D-F Group-level (all mice) comparisons for the original dataset: field variance (D), coupling variance (E), and coupling asymmetry (F). G-I: Plots analogous to D-F for shuffle-subtracted parameter variances and coupling asymmetry. Each subplot of D-I contains the p-values of Wilcoxon signed-rank tests for the active vs. passive conditions.
We assessed key features identified in the exemplary mouse (Fig. 6) across all mice by comparing them with trial-shuffled data.
First, the firing rate profiles with reduced activity in the active conditions found in Fig. 6A were consistently observed across all mice with a few exceptions (Supplementary Fig. S2 and S3). We compared the mean and sparsity of the firing rate distributions of individual neurons between the two conditions across all 37 mice (Supplementary Fig. S4). Sparsity of a non-negative firing rate distribution was quantified by the coefficient of variation (CV) [63]. The V1 neurons exhibited diminished and sparser firing rate distributions in the active condition than in the passive condition, as confirmed by the reduced mean spike rates ($p=1.556\times 10^{-8}$, Wilcoxon signed-rank test) and increased CVs ($p=8.35\times 10^{-8}$, Wilcoxon signed-rank test).
@@
Next, we assessed key statistical features of the estimated parameters of the state-space kinetic Ising model. Figure 7A-C illustrates these features for an exemplary mouse ($574078$). Figure 7A, B shows distributions of time-averaged fields $\theta_{i,t}$ and couplings $\theta_{ij,t}$ under the active and passive conditions, while Figure 7C shows a scatter plot of time-averaged reciprocal couplings $\theta_{ij,t}$ vs $\theta_{ji,t}$ to evaluate coupling asymmetry. In the active condition, the medians of field parameters decreased, reflecting reduced firing rates, while the medians of couplings remained near zero in both conditions. Field and coupling parameter variances increased, and coupling asymmetry strengthened in the active condition. These trends were consistent across all mice (Fig. 7D-F). These characteristics represent key aspects of neural dynamics that are closely related to entropy flow, although they are not entirely independent of each other.
@@
While increased parameter variabilities and coupling asymmetry were observed under the active condition, they may be influenced by the lower neuronal activity. To examine this, we compared results with trial-shuffled data across all mice. Figures 7G-I show field and coupling variances in both conditions, adjusted by subtracting shuffled data values for each mouse. Notably, observed values in both active and passive conditions were significantly higher than shuffled data: $p=2.91\times 10^{-11}$ (active), $p=1.103\times 10^{-7}$ (passive) for fields, $p=4.676\times 10^{-8}$ (active), $p=1.455\times 10^{-11}$ (passive) for couplings (Wilcoxon signed-rank test). Note that the observed significant heterogeneity in the field parameters is likely associated with the coupling heterogeneity. These results confirm that the variability observed in active or passive conditions is not explained by noise couplings. The coupling asymmetry was higher than shuffled results only for the active condition ($p=1.185\times 10^{-5}$(active) and $p=0.1287$ (passive) for asymmetry).
@@
Comparisons of these significant changes of the parameter variability (i.e., shuffled results subtracted) between the active and passive conditions showed significantly greater values in the active condition ($p=8.273\times 10^{-4}$ for fields, $p=6.421\times 10^{-4}$ for couplings, Wilcoxon signed-rank test, Fig.7G,H), indicating greater variabilities in both field and coupling parameters during active behavior. A similar analysis of the mean couplings across mice revealed slightly but significantly larger values under the active condition (Supplementary Fig. S5). In contrast, coupling asymmetry showed no significant difference ($p=0.1287$, Wilcoxon signed-rank test, Fig.7I). The lack of statistically discernible change in asymmetry in the effective couplings accords with the use of the proposed mean-field method for comparing the coupling effect, which primarily arises from variability change. These findings validate enhanced parameter variability in the sparse neuronal activity during active engagement.
Mouse V1 neurons: Entropy flow dynamics
@@
Figure 8: Estimated entropy flow dynamics. A Time courses of entropy flow for the original (solid lines) and shuffled data (dash-dot lines) under the active (red) and passive (blue) conditions. The dashed lines show the corresponding population-averaged spike rates. B Total entropy flows summed across all time bins for each mouse under active and passive conditions. Data from the same mouse are connected by a line. C Shuffle-subtracted total entropy flow (original - shuffle), shown for each mouse under the active and passive conditions.
@@
Using the estimated parameters of the state-space kinetic Ising model, we computed entropy flow dynamics. Figure 8A shows the time-varying entropy flow of a representative mouse ($574078$) under the active and passive conditions (red and blue solid lines, respectively). In both cases, transient increases in entropy flow coincided with declines in the mean population spike rate (dashed lines). Similar patterns appeared across all mice analyzed (Supplementary Fig. S6). These increases align with the second law, indicating that greater entropy dissipation is required when the system is transitioning to a lower entropy state, characterized by reduced firing rates.
@@
The entropy flow time courses for this mouse showed no clear differences between the active and passive conditions. To assess population-level effects, we analyzed all 37 mice and computed total entropy flow across time bins for each condition (Fig. 8B). The comparison revealed significantly lower total entropy flow in the active condition ($p=0.01159$, Wilcoxon signed-rank test). Note that neurons exhibited reduced firing rates (Supplementary Fig. S3) and increased parameter variability (Fig. 7D,E) during the active condition.
@@
To isolate the effect of couplings, we compared the observed total entropy flows with shuffled data results (Fig.8C). The estimated entropy flow for shuffled data includes the impact of firing rate dynamics and estimation error on couplings from other neurons; therefore, subtracting shuffling results from observed entropy flow isolates contributions of couplings among different neurons beyond the sampling fluctuation. Positive values of the shuffle-subtracted total entropy flow in both conditions indicate that the couplings caused a significant entropy flow increase ($p=1.455\times 10^{-11}$ for active, $p=1.455\times 10^{-11}$ for passive, Wilcoxon signed-rank test). These shuffle-subtracted entropy flows behave in agreement with the theoretical prediction by the Sherrington-Kirkpatrick model [37]. In the active condition, the increased coupling variability (and asymmetry) from the shuffle-subtracted values were positively correlated with the shuffle-subtracted entropy flows, while the increased field heterogeneity was negatively correlated (Supplementary Fig. S7A-C). These effects disappeared in the passive condition, possibly due to small changes in the variabilities and asymmetry introduced by shuffling (Supplementary Fig. S7D-F).
@@
We analyzed the differences in coupling-related entropy flows between the active and passive conditions for all mice (Fig.8C). The result shows no significant difference between the two conditions ($p=0.1448$, Wilcoxon signed-rank test). However, coupling-related entropy flows of indistinguishable magnitude emerged under distinct neural activity states: sparser, lower activity with increased variability in field and coupling parameters in the active condition; and less sparse, higher activity with reduced variability in the passive condition. Thus, coupled with the previous results, this result indicates that the greater coupling variability in the active condition led to increased total entropy flow, making it comparable to the passive condition despite significantly sparser firing rate distributions. Consistent with this view, a recent study by Aguilera et al. using this dataset complementarily reported that a lower bound on entropy production, derived under steady-state assumptions using a variational framework, was higher in the active condition when normalized per spike [13].
Mouse V1 neurons: Model-based perturbation analysis
@@
Figure 9: Model-based perturbation analysis A Entropy flow $\sigma_{t}^{\rm{flow}}$ of mouse 574078 in the active condition, computed after rescaling the fitted parameters as $\boldsymbol{\theta}\to\beta\boldsymbol{\theta}$. The dashed line indicates $\beta=1$. B Forward entropy flow, $\sigma_{t}^{\mathrm{forward}}$. C Backward entropy flow, $\sigma_{t}^{\mathrm{backward}}$. D Shuffle-subtracted entropy flow in the active condition, isolating the contribution of interactions beyond firing rate dynamics. Entropy flow driven by interactions peaks at $\beta<1$. E Shuffle-subtracted entropy flow in the passive condition. F Comparison of shuffle-subtracted entropy flow between active and passive conditions across all mice, showing significantly higher values in the active condition (Wilcoxon test, $p=1.455\times 10^{-10}$). Shuffle-subtracted entropy flow is obtained over a low-gain range $\beta\in[0.2,1.0]$ and across all bins. Lines connect active (left) to passive (right) for each mouse.
@@
To further elucidate the difference in the estimated entropy flow in active and passive conditions, we performed a model-based perturbation analysis by rescaling the fitted model parameters as $\boldsymbol{\theta}\to\beta\boldsymbol{\theta}$ and computing the resulting entropy flow to assess its sensitivity to parameter perturbation. An example result from mouse $574078$ (Fig.9A) shows that the entropy flows during stimulus presentation and waiting (gray image) periods exhibited distinct behaviors in response to the rescaling. The transient increases in entropy flow caused by firing rate reduction after stimulus onset and offset persisted as the scaling parameter $\beta$ increased. In contrast, we observed that the entropy flow peaked at $\beta<1$ during the waiting period, where the neural activity is relatively stationary.
@@
Both forward and reverse conditional entropies ($\sigma_{t}^{\rm forward}$ and $\sigma_{t}^{\rm backward}$ in Eq. 12) decreased with increasing $\beta$ during the waiting period (Fig.9B,C), indicating that both processes became more deterministic. This trend suggests that, as $\beta$ increases, the system transitions from a disordered phase toward a ferromagnetic phase, rather than into a quasi-chaotic regime [43]. Thus, these results indicate that the subsampled neural population during this period operates in a subcritical regime.
@@
By subtracting the entropy flow estimated from trial-shuffled data, which preserves only firing rate dynamics, we confirmed that the two bands of increased entropy flow associated with stimulus presentation are attributable to firing rate changes, whereas the increase at $\beta<1$ arises from interactions, since the former disappeared but the latter persisted after shuffle subtraction (Fig.9D). The interaction-driven entropy flow revealed by parameter scaling was stronger during the active condition than the passive condition (Fig.9D,E), a result confirmed across all mice (Fig.9F). Notably, the previous analysis at $\beta=1$ showed no difference in shuffle-subtracted (i.e., interaction-driven) entropy flow between the two conditions (Fig.8C). Thus, the model-based perturbation analysis uncovered differences in entropy flow between active and passive states that were not apparent at $\beta=1$.
Mouse V1 neurons: Entropy flow and behavioral performance
@@
Figure 10: Correlations between neural dynamics and behavioral performance. A A histogram of firing rate differences (active-passive) across neurons for mouse 574078 with an image im036_r. A skewness was used as a sparsification index. B The skewness vs behavioral performance (mean d-prime) for 37 mice with three images. C Mean-field rates of individual neurons (colored circles) as a function of time-average mean $g_{i}$ and variability $\Delta_{i}$ of their inputs for mouse 574078. The background color indicates the theoretical mean-field rate under the steady-state (Eq. 19). D Mean-field entropy flow of individual neurons (colored circles) as a function of time-average mean and variability of their inputs (mouse 574078). The background color shows theoretical entropy flow under the steady-state (Eq. 20). E Entropy flow difference (active - passive) normalized by activity rate vs behavioral performance for 37 mice with three images. F Entropy flow difference normalized by activity rate obtained from trial-shuffled data vs behavioral performance. In E and F, there is one outlier mouse below 20-20 in the ordinate, which was included in the statistical analysis.
@@
Finally, we investigated the relationship between neural dynamics and behavioral performance across individual mice. We quantified task performance by the sensitivity index $d^{\prime}$ (mean d-prime) defined as the difference between the z-transformed hit and false-alarm rates (Supplementary Note 4). In the following analyses, we extended the analysis to include two additional images (im012_r and im115_r).
@@
First, we examined how sparseness, assessed from individual neurons’ activity rates, relates to behavioral outcomes. As shown in Supplementary Fig. S3, neuronal activities were significantly reduced under active conditions, accompanied by increased sparsity of firing rate distributions. To further characterize this effect, we examined whether the reduction was uniform across neurons or driven by a subset of neurons by computing the skewness of the firing rate difference between active and passive conditions (Fig. 10A). A uniform reduction results in a skewness of zero, whereas negative skewness indicates that only some neurons decreased their activity, reflecting the sparsification. We found that this sparsification index was significantly correlated with behavioral performance measured by the d-prime, indicating that task engagement is reflected in changes in sparsity quantified at the level of individual neurons’ activity rates (Fig. 10B).
@@
Having established the link between activity-rate sparsity and behavior, we next turned to entropy flow to ask whether it provides additional explanatory power beyond rate changes alone. The variability of effective couplings was significantly higher during the active condition. To gain insight into the contributions of couplings to entropy flow, we computed the activity rate and mean-field entropy flow of individual neurons as a function of the mean and variability of their inputs (Fig. 10C,D). Theoretically, in the low-input and stationary regime, entropy flow increases with both higher mean input and greater variability (Eq. 20, background color in Fig. 10D). We observed that neurons receiving high mean input tended to have less variable inputs, whereas neurons with low mean input exhibited larger variability (colored circles). These results suggest that total entropy flow is shaped not only by high-input (typically high-firing) neurons but also by low-input neurons with high variability.
@@
These patterns imply two sources of entropy flow: (i) mean-input-driven contributions that track high firing, and (ii) variability-driven contributions that can be substantial even at low firing. To focus more on the latter, we considered entropy flow per activity rate. This normalization reduces the direct dependence on mean rate and makes variability-driven effects, particularly those arising in low-rate neurons, observable on equal footing with high-rate effects. The shift in mean entropy flow per activity rate across individual neurons (active - passive) was significantly correlated with behavioral performance (Fig. 10E). Moreover, this correlation was weaker and non-significant for trial-shuffled data, indicating contributions from highly variable couplings during active conditions (Fig. 10F). This finding suggests that the thermodynamic cost per spiking activity is related to mouse performance, with couplings contributing in addition to activity-rate sparsity.
@@
As an alternative explanation, behavioral performance could be related to entropy-flow changes concentrated in high-firing neurons. We therefore tested whether neurons with higher spike rates tended to increase entropy flow during active engagement in mice with higher task performance (Supplementary Note 5, Supplementary Fig. S8). While this tendency correlated significantly with behavioral performance for one image, it was not significant for the other two images. We therefore infer that high-firing-based changes alone cannot consistently account for performance differences. Instead, the more robust association with entropy flow per activity rate supports a complementary role of variability-driven, coupling-mediated contributions – including those from low-rate neurons – in explaining behavioral performance.